how much-distilled water should a nurse add to reduce 20 oz of a 25% antiseptic solution to a solution that is 10% antiseptic.
Answer by mananth(16946)


-------------- percent ---------------- Amount
Solution I 25 ---------------- 20 oz
Water 0 ---------------- x oz
Mixture 10 ---------------- 20 + x oz

25 * 20 + 0 x = 10 ( 20 + x )
500 + 0 x = 200 + 10 x
0 x -10 x = -500 + 200
-10 x = -300
/ -10
x = 30 oz Water

m.ananth@hotmail.ca